# Simple homotopy types of even dimensional manifolds

### v2

@@ 2.4. The Whitehead torsion of a homotopy equivalence
Now let $X$ and $Y$ be finite CW complexes with universal covers $\widetilde{X}$ and $\widetilde{Y}$, and let $F:=\pi_{1}(X)$ and $G:=\pi_{1}(Y)$. The cellular chain complex of $Y$ with $\mathbb{Z}G$ coefficients is $C_{*}(Y;\mathbb{Z}G)=\mathbb{Z}G\otimes_{\mathbb{Z}G}C_{*}(\widetilde{Y})\cong C_{*}(\widetilde{Y})$, which is a finitely generated, free, left $\mathbb{Z}G$-module chain complex. Choose a lift of each cell of $Y$ in $\widetilde{Y}$ to obtain a basis of $C_{*}(\widetilde{Y})$, which is well-defined up to ordering, signs and multiplication by elements of $G$. Similarly, the cellular chain complex of $X$ with $\mathbb{Z}F$ coefficients, $C_{*}(X;\mathbb{Z}F)\cong C_{*}(\widetilde{X})$, is a finitely generated, free, left $\mathbb{Z}F$-module chain complex with a basis well-defined up to ordering and multiplication by elements of ±$\pm F$.
@@
so this is also a finitely generated, free chain complex with a basis that is well-defined up to ordering and multiplication by elements of ±$\pm G$.
@@
The Whitehead torsion of the cellular homotopy equivalence $f\colon X\to Y$ is $\tau(f):=\tau(f_{*})$, where $f_{*}\colon C_{*}(X;\mathbb{Z}G^{\theta})\rightarrow C_{*}(Y;\mathbb{Z}G)$ is the induced chain homotopy equivalence.
By Remark 2.14, it follows that $\tau(f_{*})$ is well-defined, even though the bases of $C_{*}(X;\mathbb{Z}G^{\theta})$ and $C_{*}(Y;\mathbb{Z}G)$ are well-defined only up to ordering and multiplication by elements of ±$\pm G$.
@@
Now we can extend the definition of Whitehead torsion to arbitrary homotopy equivalences. If $f\colon X\to Y$ is a homotopy equivalence between finite CW complexes, then it is homotopic to a cellular homotopy equivalence $f^{\prime}\colon X\to Y$, and we define $\tau(f)\coloneqq\tau(f^{\prime})$. By Proposition 2.29 this is independent of the choice of $f^{\prime}$. Moreover, it follows that if $f,g\colon X\to Y$ are arbitrary homotopy equivalences and $f\simeq g$, then $\tau(f)=\tau(g)$.
Now that we have defined Whitehead torsion, it is worth recalling its key rôle: Theorem 2.2 states that a homotopy equivalence $f\colon X\to Y$ between CW complexes $X$ and $Y$ is simple if and only if its Whitehead torsion $\tau(f)=0\in\operatorname{Wh}(\pi_{1}(Y))$.
We collect a few key properties of Whitehead torsion.
@@
Let $X$, $Y$, and $Z$ be finite CW complexes and let $f\colon X\to Y$ and $g\colon Y\to Z$ be homotopy equivalences. Then
@@
The following generalises an observation of Wall. He pointed this out in the case where $M$, $N$ are simple Poincaré complexes [Wa74, p. 612], i.e. finite Poincaré complexes $X$ for which the chain duality isomorphism $C^{*}(\widetilde{X})\to C_{*}(\widetilde{X})$ is a simple chain homotopy equivalence. At the time, it was known that smooth manifolds are simple Poincaré complexes [Wall-SCM, Theorem 2.1] but the case of topological manifolds was not established until the work of Kirby-Siebenmann [Kirby-Siebenmann:1977-1, III.5.13].
Proposition 2.35.
Let $M$ and $N$ be closed $n$-dimensional topological manifolds. Let $G=\pi_{1}(N)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, and let $f\colon M\to N$ be a homotopy equivalence. Then $\tau(f)\in\mathcal{J}_{n}(G,w)$.
Proof.
@@
Let $F=\pi_{1}(M)$ and $\theta=\pi_{1}(f)\colon F\to G$. Since $f$ is a homotopy equivalence, $f_{*}([M])=\pm[N]$, where $[M]$ and $[N]$ denote the (twisted) fundamental classes of $M$ and $N$ respectively. Since Poincaré duality is given by taking cap product with the fundamental class, and the cap product is natural, we get a diagram of left $\mathbb{Z}G$-module chain complexes that commutes up to chain homotopy and sign:
${C^{n-*}(M;\mathbb{Z}G^{\theta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{PD}}$${C^{n-*}(N;\mathbb{Z}G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{C^{n-*}(f)}$$\scriptstyle{\operatorname{PD}}$${C_{*}(M;\mathbb{Z}G^{\theta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{C_{*}(f)}$${C_{*}(N;\mathbb{Z}G)}$
Poincaré duality is a simple chain homotopy equivalence by [Kirby-Siebenmann:1977-1, III.5.13], so we have that $\tau(\operatorname{PD}\colon C^{n-*}(N;\mathbb{Z}G)\to C_{*}(N;\mathbb{Z}G))=0$. By Lemmas 2.26 and 2.27, the Poincaré duality map
@@
$\operatorname{PD}\colon C^{n-*}(M;\mathbb{Z}G^{\theta})\cong C^{n-*}(M;\mathbb{Z}F)_{\theta^{-1}}\to C_{*}(M;\mathbb{Z}G^{\theta})\cong C_{*}(M;\mathbb{Z}F)_{\theta^{-1}}$
is simple too. It follows from Proposition 2.15 and Lemma 2.16 that $\tau(C^{n-*}(f))+\tau(C_{*}(f))=0$. Therefore it is enough to prove $\tau(C^{n-*}(f))=(-1)^{n}\overline{\tau(C_{*}(f))}$.
@@
By Lemma 2.19, we know that $\tau(C^{n-*}(f))=(-1)^{n}\tau(C^{-*}(f))$, and it follows from Definition 2.20 that $\tau(C^{-*}(f))=\tau(C^{*}(f))$. Finally Lemma 2.24 shows that $\tau(C^{*}(f))=\overline{\tau(C_{*}(f))}$. We deduce $\tau(C^{n-*}(f))=(-1)^{n}\overline{\tau(C_{*}(f))}$, completing the proof. ∎
@@
Suppose that $W$ is an $h$-cobordism between closed $n$-dimensional manifolds $M$ and $N$. Let $G=\pi_{1}(W)$ and let $w\colon G\rightarrow\left\{\pm 1\right\}$ be the orientation character of $W$. We will write $\tau(W,M)$ and $\tau(W,N)$ to denote the Whitehead torsion of the inclusions $M\rightarrow W$ and $N\rightarrow W$ respectively. We will refer to the composition $N\to M$ of the inclusion $N\rightarrow W$ and the homotopy inverse of the inclusion $M\rightarrow W$ as the homotopy equivalence induced by $W$.

### v3

@@ 2.4. The Whitehead torsion of a homotopy equivalence
[…]
@@
so this is also a finitely generated, free chain complex with a basis that is well-defined up to ordering and multiplication by elements of $\pm G$.
@@
The Whitehead torsion $\tau(f)\in\operatorname{Wh}(\pi_{1}Y)$ of the cellular homotopy equivalence $f\colon X\to Y$ is $\tau(f):=\tau(f_{*})$, where $f_{*}\colon C_{*}(X;\mathbb{Z}G^{\theta})\rightarrow C_{*}(Y;\mathbb{Z}G)$ is the induced chain homotopy equivalence.
By Remark 2.12, it follows that $\tau(f_{*})$ is well-defined, even though the bases of $C_{*}(X;\mathbb{Z}G^{\theta})$ and $C_{*}(Y;\mathbb{Z}G)$ are well-defined only up to ordering and multiplication by elements of $\pm G$.
@@
We can now extend the definition of Whitehead torsion to arbitrary homotopy equivalences. If $f\colon X\to Y$ is a homotopy equivalence between finite CW complexes, then it is homotopic to a cellular homotopy equivalence $f^{\prime}\colon X\to Y$, and we define $\tau(f):=\tau(f^{\prime})$. By Proposition 2.28 this is independent of the choice of $f^{\prime}$. Moreover, it follows that if $f,g\colon X\to Y$ are arbitrary homotopy equivalences and $f\simeq g$, then $\tau(f)=\tau(g)$.
We collect a few key properties of simple homotopy equivalences and Whitehead torsion.
Proposition 2.29.
Proof.
@@
Let $X$, $Y$ and $Z$ be finite CW complexes and let $f\colon X\to Y$ and $g\colon Y\to Z$ be homotopy equivalences. Then
@@
It will be important to have formulae for how Whitehead torsion behaves under products and in fibre bundles.
@@
Proposition 2.34 ([Anderson, Theorem A]).
For $i=1,2$, let $F_{i}\xrightarrow[]{j_{i}}E_{i}\xrightarrow[]{p_{i}}B_{i}$ be fibre bundles of finite CW complexes, and let
${F_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{1}}$$\scriptstyle{h}$${E_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{1}}$$\scriptstyle{g}$${B_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${F_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{2}}$${E_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{2}}$${B_{2}}$
be a fibre homotopy equivalence. Then there is a homomorphism $p_{2}^{*}\colon\operatorname{Wh}(\pi_{1}(B_{2}))\rightarrow\operatorname{Wh}(\pi_{1}(E_{2}))$ such that
$\tau(g)=p_{2}^{*}(\tau(f))+\chi(B_{2})\cdot{j_{2}}_{*}(\tau(h)).$
2.5. The Whitehead torsion of a homotopy equivalence between manifolds
In this section we specialise further to the case of manifolds, and see that $\mathcal{J}_{n}(G,w)$ and $\mathcal{I}_{n}(G,w)$ arise naturally.
[…]
@@ 2.5. The Whitehead torsion of a homotopy equivalence between manifolds
Let $M$ and $N$ be closed $n$-dimensional topological manifolds. Let $G=\pi_{1}(M)$ with orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, and let $f\colon N\to M$ be a homotopy equivalence. Then $\tau(f)\in\mathcal{J}_{n}(G,w)$.
@@
Let $F=\pi_{1}(N)$ with orientation character $v\colon F\rightarrow\left\{\pm 1\right\}$, and let $\theta=\pi_{1}(f)\colon F\to G$. Then $(\mathbb{Z}^{w})^{\theta}=\mathbb{Z}^{w\circ\theta}$, and since $f$ is a homotopy equivalence, it induces an isomorphism $H_{n}(N;\mathbb{Z}^{w\circ\theta})\xrightarrow{\cong}H_{n}(M;\mathbb{Z}^{w})\cong\mathbb{Z}$. It follows from Poincaré duality that $v$ is the only homomorphism $F\rightarrow\left\{\pm 1\right\}$ such that $H_{n}(N;\mathbb{Z}^{v})\cong\mathbb{Z}$; indeed, for any such homomorphism $u$, $H_{n}(N;\mathbb{Z}^{u})\cong H^{0}(N;\mathbb{Z}^{u\cdot v})$, which is $\mathbb{Z}$ if and only if $u=v$. Thus $v=w\circ\theta$.
Consider the following diagram of left $\mathbb{Z}G$-module chain complexes:
${C^{n-*}(N;\mathbb{Z}G^{w\theta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{PD}}$${C^{n-*}(M;\mathbb{Z}G^{w})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{C^{w(n-*)}(f)}$$\scriptstyle{\operatorname{PD}}$${C_{*}(N;\mathbb{Z}G^{\theta})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{C_{*}(f)}$${C_{*}(M;\mathbb{Z}G)}$ (2.1)
@@
The horizontal maps are the chain and cochain maps induced by $f$. The Poincaré duality map $\operatorname{PD}\colon C^{n-*}(M;\mathbb{Z}G^{w})\to C_{*}(M;\mathbb{Z}G)$ is given by taking cap product with a cycle representing the twisted fundamental class of $M$, a generator $M]\in H_{n}(M;\mathbb{Z}^{w})\cong\mathbb{Z}$ (using $\mathbb{Z}^{w}\otimes_{\mathbb{Z}}\mathbb{Z}G^{w}=(\mathbb{Z}G^{w})^{w}=\mathbb{Z}G$). Similarly, taking cap product with a representative of $N]\in H_{n}(N;\mathbb{Z}^{v})$ defines $\operatorname{PD}\colon C^{n-*}(N;\mathbb{Z}G^{w\theta})\to C_{*}(N;\mathbb{Z}G^{\theta})$. Here we write $\mathbb{Z}G^{w\theta}=(\mathbb{Z}G^{w})^{\theta}$ for brevity, and we have $\mathbb{Z}G^{w\theta}=\mathbb{Z}G^{\theta v}$ (by the definitions, using that $v=w\circ\theta$) and $\mathbb{Z}^{v}\otimes_{\mathbb{Z}}\mathbb{Z}G^{\theta v}=\mathbb{Z}G^{\theta}$. Since $f$ is a homotopy equivalence, $f_{*}([N])=\pm[M]$, so by the naturality of the cap product, diagram (2.1) commutes up to chain homotopy and sign.
Note that by Lemma 2.25 there are isomorphisms
$C_{i}(N;\mathbb{Z}G^{\theta})=\mathbb{Z}G^{\theta}\otimes_{\mathbb{Z}F}C_{i}(\widetilde{N})\cong C_{i}(\widetilde{N})_{\theta^{-1}}=C_{i}(N;\mathbb{Z}F)_{\theta^{-1}}$ (2.2)
$C^{i}(N;\mathbb{Z}G^{w\theta})$ $=C^{i}(N;\mathbb{Z}G^{\theta v})=\operatorname{Hom}_{\mathbb{Z}F}^{\mathrm{lr}}(C_{i}(\widetilde{N}),\mathbb{Z}G^{\theta v})=\operatorname{Hom}_{\mathbb{Z}F,v}^{\mathrm{lr}}(C_{i}(\widetilde{N}),\mathbb{Z}G^{\theta})$
@@
$\begin{cases}\cong\operatorname{Hom}_{\mathbb{Z}F,v}^{\mathrm{lr}}(C_{i}(\widetilde{N}),\mathbb{Z}F)_{\theta^{-1}}=\operatorname{Hom}_{\mathbb{Z}F}^{\mathrm{lr}}(C_{i}(\widetilde{N}),\mathbb{Z}F^{v})_{\theta^{-1}}=C^{i}(N;\mathbb{Z}F^{v})_{\theta^{-1}}\\ =\operatorname{Hom}_{\mathbb{Z}G,w}^{\mathrm{lr}}(C_{i}(\widetilde{N})_{\theta^{-1}},\mathbb{Z}G)=(C_{i}(\widetilde{N})_{\theta^{-1}})^{w*}\cong C_{i}(N;\mathbb{Z}G^{\theta})^{w*}\end{cases}$ (2.3)
We also have, by the definitions, that
$C^{i}(M;\mathbb{Z}G^{w})=\operatorname{Hom}_{\mathbb{Z}G}^{\mathrm{lr}}(C_{i}(\widetilde{M}),\mathbb{Z}G^{w})=C_{i}(\widetilde{M})^{w*}=C_{i}(M;\mathbb{Z}G)^{w*}.$
Using this and the second line of (2.3), $C^{w*}(f)$ can be identified with the $\vphantom{C}{}^{w*}$-dual of $C_{*}(f)$, i.e. $C^{wi}(f)=C_{i}(f)^{w*}$ for all $i$.
Poincaré duality is a simple chain homotopy equivalence by [Kirby-Siebenmann:1977-1, III.5.13], so we have that $\tau(\operatorname{PD}\colon C^{n-*}(M;\mathbb{Z}G^{w})\to C_{*}(M;\mathbb{Z}G))=0$. By the first line of (2.3), we have $C^{n-*}(N;\mathbb{Z}G^{w\theta})\cong C^{n-*}(N;\mathbb{Z}F^{v})_{\theta^{-1}}$, and by (2.2) $C_{*}(N;\mathbb{Z}G^{\theta})\cong C_{*}(N;\mathbb{Z}F)_{\theta^{-1}}$. Using this, we can identify the Poincaré duality map
$\operatorname{PD}\colon C^{n-*}(N;\mathbb{Z}G^{w\theta})\to C_{*}(N;\mathbb{Z}G^{\theta})$
with
$\operatorname{PD}\colon C^{n-*}(N;\mathbb{Z}F^{v})_{\theta^{-1}}\to C_{*}(N;\mathbb{Z}F)_{\theta^{-1}}.$
By Lemma 2.26, the Whitehead torsion of the latter map is given by $\theta_{*}$ applied to the Whitehead torsion of $\operatorname{PD}\colon C^{n-*}(N;\mathbb{Z}F^{v})\to C_{*}(N;\mathbb{Z}F)$. But this torsion vanishes, again by [Kirby-Siebenmann:1977-1, III.5.13]. Thus all of the above Poincaré duality maps are simple. It then follows from (2.1), Proposition 2.13, and Lemma 2.14 that $\tau(C^{w(n-*)}(f))+\tau(C_{*}(f))=0$. Therefore it is enough to prove that $\tau(C^{w(n-*)}(f))=(-1)^{n}\overline{\tau(C_{*}(f))}$.
@@
By Lemma 2.16, we know that $\tau(C^{w(n-*)}(f))=(-1)^{n}\tau(C^{w(-*)}(f))$, and by Definition 2.17 $\tau(C^{w(-*)}(f))=\tau(C^{w*}(f))$. Finally Lemma 2.23 shows (the second equality of) $\tau(C^{w*}(f))=\tau(C_{*}(f)^{w*})=\overline{\tau(C_{*}(f))}$. We deduce that $\tau(C^{w(n-*)}(f))=(-1)^{n}\overline{\tau(C_{*}(f))}$, completing the proof. ∎
@@
Suppose that $W$ is an $h$-cobordism between closed $n$-dimensional manifolds $M$ and $N$. Let $G=\pi_{1}(W)$ and let $w\colon G\rightarrow\left\{\pm 1\right\}$ be the orientation character of $W$. We write $\tau(W,M)$ and $\tau(W,N)$ to denote the Whitehead torsion of the inclusions $M\rightarrow W$ and $N\rightarrow W$ respectively. We refer to the composition $N\to M$ of the inclusion $N\rightarrow W$ and the homotopy inverse of the inclusion $M\rightarrow W$ as the homotopy equivalence induced by $W$.